OpenStudy (anonymous):

add or subract simplify if possible (4a+5b)/2 - (4a - 5b)/2

OpenStudy (anonymous):

Since they have a common denominator you "add/subtract" the tops and place that answer over the denominator " the 2". When you have a subtraction problem, you need to change the - to a + and change the sign of each term in the parentheses after the minus sign... so you have (4a+5b)/2 + (-4a + 5b)/2 I change the 4a to a -4a and the -5b to a +5b Then you add like terms... 4a - 4a = 0 (the a's are gone. 5b + 5b = 10b So you have 10b/2 = 5b

OpenStudy (anonymous):

thanks a bunch :)